OpenStudy (anonymous):

Solve 152x = 36. Round to the nearest ten-thousandth

OpenStudy (anonymous):

i am going to make a guess that this is \[15\times 2^x=36\] let me know if that is right or if i am mistaken

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then first step is to divide by 15 to get \[2^x=\frac{36}{15}=\frac{12}{5}\]

OpenStudy (jiteshmeghwal9):

now i think u must change this into logarithmic form :)

OpenStudy (anonymous):

then to solve \[v^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] or in your case \[x=\frac{\ln(2.4)}{\ln(2)}\]

OpenStudy (anonymous):

typo there, should be \[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] in english "the log of the total divided by the log of the base "

OpenStudy (anonymous):

since you are going to have to use a calculator it is easier to write \(\frac{12}{5}=2.4\)

OpenStudy (jiteshmeghwal9):

\[\LARGE{\log_{2}{12/5}=x}\]

OpenStudy (jiteshmeghwal9):

isn't it an easy way @satellite73

OpenStudy (anonymous):

@jiteshmeghwal9 it is the same answer, unfortunately you will not get a decimal approximation from this since you do not have log base 2 on your calculator if you want to know the number as decimal you h ave to use the change of base formula for either base e or base 10 because those are the ones available to you

OpenStudy (jiteshmeghwal9):

k ! thanx @satellite73 :)